### v1

@@ 4.2.2 Results
As suggested by the convergence analysis, the effect of increasing model evaluations varied by method. For CES and EKS, mean AEs decreased by approximately 50% from 200 to 5400 evaluations, and these differences were statistically significant (Table SM1), Figure 4). Similarly, BOED1000 achieved a 50% reduction in mean AE compared to BOED200, although this difference was not significant. In contrast, increasing model evaluations for HM and GBOED had little effect on AEs. Interestingly, GBOED200 achieved a slightly lower mean AE than GBOED1000 (Figure 4), reinforcing our earlier observation that additional model evaluations do not necessarily improve calibration performance for GBOED.
Out of the methods utilizing fewer evaluations (200), GBOED200 achieved the lowest mean AE and performed comparably to EKS5400. Among the methods with higher evaluation counts, EKS5400 achieved the lowest mean AE (Figure 4). All pairwise comparisons from the linear AE model are provided in SM2, SM1.
Results for score were largely consistent with those for AE, with some notable differences. Like results for AE, LHS200 and LHS1000 had significantly worse (lower) scores compared to all other methods (Table 1, Figure 4). Notably, LHS-based methods were the only ones with extreme score outliers (-100 or less). This finding emphasizes the importance of an initial calibration step to achieve robust UQ. EKS5400 again achieved the highest mean score, significantly outperforming six methods: LHS200, LHS1000, BOED200, BOED1000, EKS200, and CES200. GBOED1000 outperformed five methods, while CES5400 and GBOED200 ranked similarly, outperforming four methods each (Table 1). All pairwise comparisons based on the linear model for score can be found in SM2, Table SM2.
Unlike the results for AE, increasing the number of model evaluations for GBOED slightly improved UQ. Although BOED methods generally underperformed, they achieved significantly better scores than LHS-based methods. Finally, while HM methods had fair performance in parameter estimation accuracy, significantly outperforming four methods, their performance in UQ was limited to outperforming only the LHS-based methods (Table 1).
Table 1: Methods in the leftmost column have significantly lower (better) mean AE values than those in the middle column (AE) and significantly higher (better) scores than those in the rightmost column (Score) at $\alpha=0.05$. For example, the estimated marginal mean AE of EKS200 in the leftmost column is significantly lower than mean AEs of BOED200, LHS200, and LHS1000 in the middle column. Significant differences are based on pairwise comparisons from the linear models for AE and score (Eq. 12).
@@ 5 Discussion and conclusions
In this study, we compared four Bayesian calibration methods (BOED, GBOED, HM, CES) along with EKS (the “Calibration” step of CES) and naive LHS-based methods for calibrating the Lorenz ’96 multiscale system. Our study provides a rigorous comparison of traditional methods for solving inverse problems (CES, HM) and less commonly used methods like BOED and GBOED. Each method was tested with both a limited number (200) and a larger number (1000–5400) of model evaluations. Additionally, we conducted a convergence analysis to understand how performance metrics evolved as the number of model evaluations increased.
Our results highlight several key insights. Traditional methods, particularly CES and its built-in calibration method EKS, performed exceptionally well in calibrating the Lorenz ’96 system. EKS slightly outperformed CES in both accuracy and UQ when using a large number (5400) of model evaluations. GBOED exhibited comparable performance to EKS in accuracy and uncertainty quantification. BOED, however, generally underperformed relative to GBOED, CES, HM, and EKS, suggesting that its design may not be well-suited to the emulator-based calibration problem as implemented here. This may be because an emulator like a GP does not generalize well far from training data so it is important to have many samples in calibration-relevant regions, which motivates the use of goal-oriented methods. HM performed reasonably well and could serve as an initial refinement of the parameter space during calibration. With a limited set of model evaluations, HM can significantly refine the parameter space and could be used as a precursor to other methods such as CES. Such an approach is especially advantageous when computational resources are limited and information about the underlying system is unknown, as it avoids wasting evaluations on implausible regions.
Our paper is one of the first to apply BOED-based methods to climate model calibration, following the earlier study by [12]. In [12] BOED was integrated into CES to guide design choices for improving seasonal time-averaged statistics through calibration. In our implementation of BOED and GBOED, the “calibration” step is already embedded within the BOED and GBOED algorithms. An interesting extension of this work could involve integrating GBOED (which outperformed BOED) into CES, potentially leveraging the strengths of both methods. The underwhelming performance of BOED may also stem from our test system, the Lorenz ’96 system, such as its relatively small parameter space or the simplicity of the problem. Alternatively, BOED’s performance may depend on choices in its implementation, such as prior distributions, emulator fitting, or the number of evaluations used for KL divergence calculations. These aspects warrant further investigation.
While GBOED demonstrated excellent performance with as few as 200 model evaluations, it came at the expense of significant computational costs. For instance, running GBOED for 1000 evaluations required approximately a week to complete on an HPC system. This high computational burden is primarily due to the repeated evaluation of the KL divergence, as well as the intermediate sampling steps required to optimize the Expected Information Gain (EIG). The additional complexity arises from the nested optimization loops and the need to repeatedly fit and query the GP emulator. Additionally, we observed mixed results for GBOED’s performance as model evaluations increased. Our comprehensive performance assessment showed that while its accuracy slightly decreased, its UQ marginally improved when model evaluations increased from 200 to 1000. This suggests that GBOED may benefit from tuning specific elements of its setup, such as the choice of priors or emulator fitting. Given its high computational cost, we do not recommend using GBOED for more than 200 evaluations unless specific computational resources or improvements in implementation efficiency are available.
Our findings also differ somewhat from those of [7], who showed that CES better characterized parameter uncertainty for the Lorenz ’96 system compared to EKS. In contrast, we found that EKS outperformed CES in both accuracy and UQ for the same system. While we implemented CES based on methods described in [7], differences in sampling or emulation steps may explain the discrepancy. For larger, more computationally intensive problems, performing as many evaluations as we did for EKS (e.g., 5400) is generally not feasible. In such cases, CES would offer a significant advantage due to its efficient integration of the calibration, emulation, and sampling steps. For smaller problems, however, EKS appears to be sufficient and is simpler to implement because it requires neither emulator fitting nor MCMC sampling.
Overall, this study provides valuable insights into the strengths and limitations of various Bayesian calibration methods. While traditional methods like EKS and CES perform reliably, GBOED shows promise as a non-traditional OED approach for calibration, it requires further optimization to manage its computational cost. BOED, while underperforming here, may still hold potential in other contexts with different implementation choices or problem complexities. Finally, while HM exhibited fair performance on its own, it could be used as a precursor to other methods. Future work should explore the performance of these methods on more complex systems, as well as investigate how implementation details affect results.

### v2

@@ Results
The effect of increasing model evaluations varied by method. For EKS, mean AE decreased by approximately 50% from 200 to 5400 evaluations, and this difference was statistically significant (Table SM1, Figure 4, top panel). For BOED and CES, mean AEs decreased with increasing evaluation count (from 0.85 to 0.72 for BOED and from 0.50 to 0.29 for CES), although these reductions were not statistically significant. Increasing model evaluations for HM and GBOED resulted in marginal decreases in AEs (Figure 4, top panel).
Out of the methods utilizing fewer evaluations (200), GBOED200, CES200, and HM200 performed best. Mean AEs of these methods were very similar, and although they were larger than those of the best-performing methods (e.g., CES5400, EKS5400, GBOED1000), the differences were not statistically significant. Among the methods with higher evaluation counts, CES5400 achieved the lowest mean AE (Figure 4, top panel). Another notable finding was that for fewer (200) model evaluations, CES resulted in considerable improvement over EKS. However, for more model evaluations, improvements were marginal (Figure 4, top panel). All pairwise comparisons from the linear AE model are provided in Table SM1.
Results for score were largely consistent with those for AE, with some notable differences. Like results for AE, LHS200 and LHS1000 had significantly worse (lower) scores compared to all other methods (Table 1, Figure 4, bottom panel). LHS-based methods were the only approaches exhibiting large, negative outliers (-50 or less). This finding emphasizes the importance of an initial calibration step to achieve robust UQ. CES5400 again achieved the highest mean score, significantly outperforming five methods: LHS200, LHS1000, BOED200, BOED1000, and EKS200. However, EKS5400, GBOED200, and GBOED1000 also performed very well and outperformed the same five methods as CES5400 (Table 1, Figure 4). HM200 and HM1000 had nearly identical mean scores and performed only slightly worse than top performing methods. All pairwise comparisons from the linear score model are provided in Table SM2. As an additional consistency check, we also evaluated the joint posterior density at the true parameter values for each of the replicates using a multivariate kernel density estimate; results are broadly consistent with those obtained using AE and score (see SM7).
Table 1: Methods in the leftmost column have significantly lower (better) mean AE values than those in the middle column (AE) and significantly higher (better) scores than those in the rightmost column (Score) at $\alpha=0.05$. For example, the estimated marginal mean AE of EKS200 in the leftmost column is significantly lower than mean AEs LHS200 and LHS1000 in the middle column. Significant differences are based on pairwise comparisons from the linear models for AE and score (Eq. 12).
@@
4 Two-layer quasi-geostrophic (QG) system
For our second example, we consider the quasi-geostrophic system (QG) [25]. It provides a good test for calibration methods, because this system is computationally demanding [66]. We use the two-layer QG model implemented in GeophysicalFlows.jl [13]. The prognostic variables are the potential vorticities (PVs) $q_{1}(x,y,t)$ and $q_{2}(x,y,t)$ in an upper and lower layer of thicknesses $H_{1}$ and $H_{2}$, respectively. Standard QG processes are included: nonlinear advection, planetary vorticity gradient ($\beta$-effect), interlayer coupling set by stratification, linear bottom drag $\mu$ in the lower layer, and small-scale dissipation via hyperviscosity $\nu$. A mean zonal shear $U$ is imposed to trigger baroclinic instability; after spin-up the flow reaches a statistically stationary eddy field.
We solve on a doubly periodic domain of size $L_{x}\times L_{y}=2\pi\times 2\pi$ using a pseudo-spectral method with $N_{x}\times N_{y}=256\times 256$ Fourier modes, 2/3 de-aliasing, and time step $\Delta t=1e^{-3}$. The model is integrated in time with a fourth-order exponential time–differencing Runge–Kutta scheme. Unless stated otherwise we non-dimensionalize by the domain scale and planetary vorticity, and use double precision. All QG runs in this paper use the same discretization and boundary conditions. The QG system has been used to test Bayesian calibration methods as in [66], but in contrast to the study in [66], we calibrate four parameters ($H_{1},H_{2},\mu,\nu$) rather than only $H_{1}$ and $H_{2}$.
4.1 Implementation details
We model observations as $y_{obs}=\mathcal{G}(\theta)+\varepsilon$, $\varepsilon\sim\mathcal{N}(0,\Gamma_{obs})$, where $\mathcal{G}(\theta)\in\mathbb{R}^{10}$ denotes the vector of ten flow statistics derived from the QG model at parameter value $\theta$ (for exact definitions see SM8). To compute $y_{obs}$, we first integrate for a spin-up of 75,000 steps to statistical stationarity. Then, the ten flow statistics are calculated over windows of length 25,000 step until 200 samples are collected. The observation vector, $y_{obs}$, is set to be the final (200th) sample. The observed covariance matrix, $\Gamma_{obs}$ is computed as the covariance over the 200 samples. This procedure to obtain $y_{obs}$ and $\Gamma_{obs}$ is thus similar to the procedure used to calculate these quantities for the Lorenz ’96 example. We calibrate
$\theta=(\mu,\nu,H_{1},H_{2}),$
with $\log$-parameters for $(\mu,\nu)$ to enforce positivity. $H_{1}$ and $H_{2}$ control the stratification and the baroclinic coupling strength (i.e., the deformation radius). They influence how energy and PV variance partition between layers and the preferred mesoscale length scales. The parameter $\mu$ is a large-scale linear damping acting primarily in the lower layer. It sets a natural damping timescale, and $\nu$ suppresses small-scale variance and extreme PV excursions. Together, $H_{1},H_{2},\mu,$ and $\nu$ shape the energy budget, scale selection, interlayer coupling, and intermittency, making them natural targets for Bayesian calibration. The $\beta$ parameter and imposed shear $U$ are fixed at $\beta=5$ and $U=[1,0]$. We consider the true parameters $\theta=(0.032,9.5\times 10^{-6},0.25,0.85)$ for $\mu,\nu,H_{1}$ and $H_{2}$. Priors are independent uniform distributions:
$\log\nu\sim\textup{Unif}(\log(5\times 10^{-6}),\log(1\times 10^{-4}))$
$\log\mu\sim\textup{Unif}(\textup{log}(0.01),\textup{log}(0.2))$
$H_{1}\sim\textup{Unif}(0.1,0.3)$
$H_{2}\sim\textup{Unif}(0.6,0.9).$
Because the QG system is computationally demanding, we calibrate it for a single parameter setting rather than across multiple replicates, as was done for the Lorenz ’96 example. We nevertheless examine the effect of computational budget by comparing results obtained using 200 and 500 QG model evaluations.
For QG calibration, we use Ensemble Kalman Inversion (EKI) [37, 64] rather than EKS. In preliminary tests, EKI produced more stable and accurate parameter estimates than EKS for moderate ensemble sizes and computational budgets. EKI has also been used successfully as the calibration step in CES for estimating parameters in general circulation models [19, 20]. The improved performance of EKI relative to EKS in this setting may be attributable to the additional stochasticity introduced by EKS combined with mild confounding between parameters $H_{2}$ and $\nu$, which can lead to biased ensemble drift. We implement EKI using the Julia package EnsembleKalmanProcesses.jl [12] with default settings.
For the 200-simulation case, we use $N=40$ ensemble members and 5 EKI iterations, while for the 500-simulation case we use $N=100$ ensemble members and 5 iterations, which aligns with other studies that use EKI for the calibration step of CES [35, 19, 20]. Statistics for ensemble members are computed in parallel using GPUs. After completion of the EKI iterations, GP emulators are trained using all ensemble inputs and corresponding outputs (the 10-dimensional vector of QG statistics), and posterior sampling is performed using ST-MCMC [6]. This MCMC procedure is applied consistently across CES and all other methods.
History Matching (HM) is performed in the same manner as for the Lorenz ’96 system, using five waves with 40 evaluations per wave for the 200-evaluation case and 100 evaluations per wave for the 500-evaluation case. BOED and GBOED are implemented analogously, using 10 and 25 batches of 20 QG evaluations, respectively, yielding total budgets of 200 and 500 evaluations. The batch size is chosen to match the HPC job limit, allowing all model evaluations within a batch to run concurrently on separate nodes. For all methods, the final posterior is obtained by training GP emulators on the full set of input–output pairs generated during calibration and applying ST-MCMC, as in the Lorenz ’96 example.
4.2 Results
We begin by examining histograms of the posterior distributions for CES, HM, BOED, and GBOED using 200 and 500 model evaluations. Distributions for $\mu$ and $\nu$ are shown in log space. For both budgets, the posterior distributions produced by all methods include the true parameter values (Figure 5, left and right panels).
Increasing the number of model evaluations from 200 to 500 generally leads to more concentrated and unimodal posteriors, most notably for BOED (blue). For parameter $\nu$, the posterior mass is more tightly concentrated around the true value for HM and BOED, whereas CES and GBOED place relatively less posterior mass near the true value for both evaluation budgets. For parameter $H_{2}$, increasing the number of model evaluations appears to slightly broaden the HM posterior, shifting mass away from the true value. Overall, increasing the number of model evaluations does not consistently result in substantial changes to posterior shape across methods, a behavior that mirrors what we observed for the Lorenz ’96 system.
Figure 5: Frequency histograms of posterior distributions of $\textup{log}\mu,\textup{log}\nu,H_{1}$ and $H_{2}$ for 200 model evaluations (left) and 500 model evaluations (right). Histograms represent 8092 values. True parameter values ($\theta=(\log 0.032,\log 9.5\times 10^{-6},0.25,0.85)$ for $\log\mu,\log\nu,H_{1}$ and $H_{2}$) are denoted by vertical black lines.
We assessed convergence by evaluating absolute error (AE) and score every 100 model evaluations, from 100 to 500 evaluations, for all methods. In contrast to the convergence behavior observed for the Lorenz ’96 multiscale system, most methods exhibited clearer improvement trends for the QG system, particularly in terms of score (Figure 6, right).
For AE, EKI showed the most pronounced improvement as the number of model evaluations increased, with substantial reductions between 100 and 500 evaluations across all parameters (Figure 6, left). CES and GBOED exhibited rapid initial improvements from 100 to 200 evaluations, followed by a plateau, suggesting diminishing returns once the calibration-relevant region was sufficiently explored. In contrast, HM and BOED achieved relatively low AE values early on and displayed comparatively little variation as additional model evaluations were added.
Score-based convergence displayed an even clearer trend. Scores for nearly all methods and parameters increased steadily with additional model evaluations, indicating improved uncertainty quantification as posterior variance decreased (Figure 6, right). Overall, these results suggest that increasing model evaluations for the QG system leads to more consistent gains in both accuracy and uncertainty quantification than observed for the Lorenz ’96 multiscale system. Overall, the increase in performance exhibited by most methods for the QG system likely reflects stronger parameter identifiability and smoother parameter–statistic relationships than in the Lorenz ’96 multiscale system. In this setting, additional model evaluations may more reliably improve emulator fidelity in calibration-relevant regions, leading to gains in uncertainty quantification and, to a lesser extent, point estimation accuracy.
Figure 6: AEs (left) and scores (right) for CES, EKI, HM, BOED, and GBOED over increasing QG model evaluations, based on true parameters $\theta=(0.032,9.5\times 10^{-6},0.25,0.85)$ for $\mu,\nu,H_{1}$ and $H_{2}$. Results for $\mu$ and $\nu$ are computed in log space.
@@ 5 Discussion and conclusions
In this study, we compared four emulator-based Bayesian calibration strategies, BOED, GBOED, History Matching (HM), and Calibrate–Emulate–Sample (CES), alongside ensemble Kalman–based approaches and baseline LHS sampling, across two chaotic dynamical systems with very different computational and structural characteristics. Our results highlight both the robustness of Bayesian calibration methods and the importance of aligning design strategies with the downstream inferential task.
For the Lorenz ’96 multiscale and quasi-geostrophic (QG) systems, CES and its calibration component EKS or EKI consistently delivered strong performance in both accuracy and uncertainty quantification. When computational budgets were limited, CES provided clear advantages over EKS alone; however, these gains diminished as the number of model evaluations increased. GBOED and HM achieved performance comparable to CES and EKS even at relatively low model evaluation counts, suggesting that explicitly targeting calibration-relevant regions of parameter space, either through goal-oriented design or implausibility-based pruning, can be as effective as integrated, ensemble-based approaches.
In contrast, standard BOED generally underperformed relative to the other methods for the Lorenz ’96 multiscale system. This behavior likely reflects a mismatch between BOED’s objective to maximize information gain about the emulator and the needs of calibration, where accuracy in posterior-relevant regions is paramount. In chaotic systems with localized regions of posterior mass, GP emulators may generalize poorly outside well-sampled regions, making globally informative designs suboptimal for inference. However, this observation provides the impetus for developing goal-oriented extensions such as our GBOED algorithm, which explicitly aligns design decisions with posterior uncertainty reduction.
Although we did not perform the same extensive replicate-based analysis for the QG system as we did for Lorenz ’96 due to computational cost, all methods produced reasonable posterior estimates even with moderate evaluation budgets. Differences between methods were less pronounced, and no single approach clearly dominated. This convergence in performance likely reflects the smoother statistical dependence of QG diagnostics on parameters, which reduces sensitivity to design choices and mitigates the risks associated with emulator extrapolation. As a result, additional model evaluations tend to provide genuinely informative refinements to the emulator in calibration-relevant regions, leading to more consistent improvements in posterior concentration and score as computational budget increases. In this setting, EKI proved more stable than EKS, possibly due to reduced stochasticity in the update step combined with mild parameter confounding.
Taken together, these results suggest that for moderately smooth, high-cost systems such as QG, multiple emulator-based calibration strategies can yield comparable performance, and method selection may be guided more by practical considerations such as implementation complexity, parallelizability, and computational robustness than by accuracy alone. For more irregular or strongly nonlinear systems, however, calibration performance benefits from concentrating computational effort in regions of parameter space that are relevant for inference (GBOED, HM), or implicitly through iterative calibration steps (CES) leading to clear advantages over standard BOED. Recent results by [62] suggest that active learning and BOED-based calibration methods yield the greatest benefit when observations strongly constrain the parameter space. This perspective helps explain why differences among CES, HM, and GBOED are relatively small for the QG system, while performance disparities are more pronounced for the Lorenz ’96 multiscale system, where the inverse problem may be less tightly constrained by the chosen summary statistics.
This work contributes to the limited literature on BOED-based calibration for chaotic and computationally intensive models [20, 62]. While standard BOED may be ill-suited to some calibration problems, goal-oriented design strategies appear promising and may be further strengthened by integration with ensemble-based frameworks such as CES. Future work should explore hybrid approaches that combine the global efficiency of ensemble methods with the targeted refinement of goal-oriented design, as well as extend these comparisons to higher-dimensional parameter spaces and alternative emulator architectures. Additionally, while batching was essential for practical deployment of BOED and GBOED on HPC systems, the choice of batch size was driven primarily by queue constraints rather than methodological considerations. The broader question of how batch size and number of iterations should be balanced under a fixed model evaluation budget and how these choices impact convergence and posterior quality remains an important topic for future study.